### v3

# New Approach to Strong Correlation: Twisting Hubbard into the Orbital Hatsugai-Kohmoto Model

Abstract
[…]
I Introduction
The paradigmatic model for strong correlations in quantum matter is that of electrons hopping on a lattice between nearest neighbors but paying the energy cost $U$ anytime they doubly occupy the same site. Naively, the energy eigenstates of this model would seem to be ordered with respect to the number of doubly occupied sites. However, this fails because the potential and kinetic terms do not commute. As a result, electrons can lower their energy by hopping to sites with single occupancy. It is for this reason that this simple model remains unsolved in any dimension exceeding unity, thereby remaining a grand challenge as it is the gold standard for Mott physics in the cuprates. The state-of-the-art remains quantum Monte CarloZheng et al. (2017); Huang et al. (2018); Qin et al. (2020, 2016); LeBlanc et al. (2015), density matrix renormalization group (DMRG)Jiang and Kivelson (2022); Zheng et al. (2017); Huang et al. (2018); Qin et al. (2020) and dynamical mean-field theory (DMFT)Georges et al. (1996); Xu et al. (2013); Deng et al. (2013) and its cluster versions such as cellular DMFTPark et al. (2008); Kancharla et al. (2008) and the dynamical cluster approximation (DCA)Maier et al. (2005); Mai et al. (2022, 2023a); Werner et al. (2009) which have all been useful in unveiling the properties of the Hubbard model including the pseudogapMaier et al. (2005); Stanescu and Kotliar (2006); Ferrero et al. (2009); Wu et al. (2018); Gull et al. (2012), superconductivityQin et al. (2020); Maier et al. (2005) and transportHuang et al. (2019); Deng et al. (2013); Brown et al. (2019) in the strange metal regime.
A natural question arises: Is there an alternative? The more tractable formulation of Mott physics is the Hatsugai-Kohmoto (HK) modelHatsugai and Kohmoto (1992) which we refer to as the band HK model in the following. It has a repulsion between any two electrons with opposite spin that doubly occupy the same momentum state. Consequently, the doubly occupied band steadily increases in energy as the repulsion increases resulting in an insulating state in a half-filled band when the interaction strength exceeds the bandwidth. In this model, double occupancy is the organizing principle for the ordering of the eigenstates. Clearly then, this is a gross simplification of the Hubbard model stemming from the locality in momentum space (long-range interactions in real space) and generating a macroscopic spin degeneracy in the singly occupied sector. In essence, the solvability of band HK rests on the commutativity of the potential and kinetic energy terms. Can such non-commutativity be put back in systematically, thereby reducing the difference with the Hubbard model, without giving up on exact solvability?
We show here that this can be done. Our work shows that two rather seemingly disparate approaches, the momentum-space or band HK model and the exact diagonalization (ED) of small Hubbard clusters share a fundamental link. Our work here is based on an orbital extension of the HK model in which each k-state is decorated by $n$ local (orbital) degrees of freedom. To show the power of the model, we present at the outset Fig. 1(a) which compares the exact Bethe ansatz results in one dimension (1d) for the ground state energy with the orbital HK (OHK) model as a function of the number of orbitals, $n$. As $n$ increases, the OHK model rapidly approaches the exact result. Already at $n=10$, the deviation from the exact result is less than $1\%$. Precisely why this agreement with the Hubbard model is expected as $n$ increases is the subject of this paper. What we establish here is that as $n\rightarrow N_{\rm sites}$, the OHK model is exactly the Hubbard model regardless of the spatial dimension. Where possible, we compare with existing methodology. In particular, we show that with a few orbitals, the OHK model recovers all known results in 1d is even more convergent that the standard tool in 1d (density matrix renormalization group (DMRG)) and is in agreement with standard results in two dimensions (2d) for the Mott transitionGeorges et al. (1996); Maier et al. (2005), and spin susceptibilityWhite et al. (1989), spectral functionSeki and Yunoki (2016); Huang et al. (2022a), double occupancyWhite et al. (1989); Xu et al. (2024), dynamical spectral weight transferMeinders et al. (1993); Eskes et al. (1991), and heat capacityDuffy and Moreo (1997); Wang et al. (2022). The nature of the rapid convergence is explained in the scaling argument in the penultimate section. In essence, the orbital HK model is an efficient and cost-effective simulator of Mott/Hubbard physics, enabling analytical insights and mumerical augmentation. This is a consequence of the flow of OHK to the Mott fixed point in which a discrete $Z_{2}$ symmetry is brokenHuang et al. (2022b).
Figure 1: Comparison of the difference between the ground-state energy with n-OHK model, $E_{\rm HK}^{n}$, and the infinite-system size Hubbard Bethe ansatz energy, $E_{\rm HB}^{\infty}$. (a) OHK is solved by ED. (b) OHK (blue dots) is solved by DMRG including as many as 40 orbitals at $U=4$. The dashed line is a machine learning polynomial regression fitting with extrapolation to $1/n\rightarrow 0$ ($n\rightarrow\infty$). The asymtotic values at $01/n=$ 0 are $-5.4\times 10^{-5}~{}(U=4)$, $-9\times 10^{-5}~{}(U=6)$, $-4\times 10^{-5}~{}(U=8)$. The fitting curves can be well represented as $f(U=4)=0.45(1/n^{1.83})$, $f(U=6)=0.51(1/n^{2.01})$ and $f(U=8)=0.45(1/n^{2.07})$.
II Local Correlations
The starting point for our analysis is Fig. 2a which shows that although the HK model has all-to-all interactions, the hole 2-point correlator, $g_{h}(r)$, (blue dots) is identical to that of the non-interacting correlator (orange squares) for a metallic density of $n=0.75$. To compensate for bifurcation of the spectral weight in the HK case, we set the density of the non-interacting system to $2\times 0.75=1.5$ and multiplied the corresponding $g_{h}(r)$ by $1/2$. See Supplemental Material (Figs. S5 and S6) for further details on the evolution of exponential decay at half-filled insulator to the algebraic fall-off in the doped metallic system. The fact that the two are coincident implies that the interacting HK and non-interacting Fermi liquid are equally local. The same is true for the 4-point and all higher correlation functions. This is significant because Wightman’s reconstruction theoremWightman (1956) ensures that if all the n-point functions are local, then so is the theory. The question arises, why does a model admit non-local interactions at the level of the Hamiltonian but exhibit purely local real-space correlations? To gain insight into this, we plot in Fig. 2b the density of states for various $U$ across the Mott transition of the half-filled band HK model, which interestingly has never been computed. Here the Mott transition to the insulating state proceeds via depletion of the density of states at the central peak at $\omega=0$, similar to that in the standard single-site DMFT simulationGeorges et al. (1996). This is telling and consistent with the locality of the n-point correlators. Ultimately, the Hamiltonian of the band HK model can be mapped onto a local model. In fact, it is identical in form to the atomic limitManning-Coe and Bradlyn (2023) of the Hubbard model with the band index ${\bf k}$ replacing the site index ${\bf i}$. A more precise statement will be made later when we complete the mapping to the full Hubbard model. In fact, we present in the conclusion a scaling argument showing how precisely the $d=\infty$ limit obtains.
Figure 2: (a) 2-point correlator showing the absence of long-range correlations in the 2d band HK model at different densities despite the all-to-all nature of the interactions in real space with $U/t=10$ and $\beta=50/t$. Here $r=r_{x}=r_{y}$. (b) The density of states under different $U$ displaying Mott transition at the half-filled band HK model with $\beta=200/t$.
III Orbital HK
Given that the band HK model is ultimately local, it should be generalizable to include the non-commutativity of the kinetic and potential energies. As pointed out previouslyManning-Coe and Bradlyn (2023), band HK can be generalized to include multiple orbitals per ${\bf k}$-state as is necessary for the adaption of this model to topological models, all of which have at least 2 atoms per unit cellMai et al. (2023b, c). Such a change produces qualitatively new physics as the hybridization between the orbitals liftsManning-Coe and Bradlyn (2023) the thermodynamic degeneracy of the band HK model while preserving only the degeneracy contraint by the symmetrySetty et al. (2023). Inspired by that, we adopt an OHK model to capture the strongly correlated dynamical mixing in a simple square lattice by redefining a set of lattice sites as orbitals. This leads to a reduction in the size of the Brillouin zone (BZ) scaling as $1/n$, referred to as the reduced Brillouin zone (rBZn), as it depends on the orbital number $n$ and shown in Fig. 3. As we will see, the convergence of this model to the exact result is faster than this, roughly scaling as $1/n^{2}$.
Figure 3: Evolution of the reduced Brillouin zone (with the constraint $N\times n=N_{\rm sites}$) as the number of orbitals ($n$) increases, leading to a purely local (in k-space) model when $n=N_{\rm sites}$.
The most general version of this model is given by
$H_{\text{OHK}}=&\sum_{{\bf k},\alpha,\alpha^{\prime},\sigma}g_{\alpha,\alpha^{\prime}}({\bf k})c_{{\bf k}\alpha\sigma}^{\dagger}c_{{\bf k}\alpha^{\prime}\sigma}-\mu\sum_{{\bf k},\alpha}n_{{\bf k}\alpha,\sigma}\\ &+\sum_{{\bf k},\alpha,\alpha^{\prime}}U_{\alpha,\alpha^{\prime}}n_{{\bf k}\alpha\uparrow}n_{{\bf k}\alpha^{\prime}\downarrow},\end{aligned}$ (1)
where $k$ is summed over the rBZn and $g(\bf k)$ is the dispersion matrix, determined by the underlying lattice hopping process and orbital setup, $\alpha(\alpha^{\prime})$ is the orbital index (ranging $1$ to $n$) and $\sigma$ is the spin. Note while the kinetic and potential energies no longer commute, the Hamiltonian still maintains the form $\sum_{k}h_{k}$ and hence is still exactly solvable and thus governed by the HK fixed pointZhao et al. (2023a); Huang et al. (2022b). Fig. 3 depicts the reduction of the Brillouin zone when $n$ orbitals decorate each k-point (see supplemental Fig. S7 for real-space counterparts). As a reduction of the Brillouin zone, in the limit of $n\rightarrow N_{\rm sites}$, the k-summation vanishes reducing the interaction term entirely to,
[…]
which is just a summation over the local degrees of freedom as in the generalized Hubbard model. The on-site Hubbard model results by just contracting the $\alpha$ and $\alpha^{\prime}$ indices. Hence, we are guaranteed to obtain Hubbard physics for $n$ sufficiently large. Rather than the $1/n$ convergence , the real convergence as computed from the difference between the computed OHK ground state energy and the exact Bethe ansatz result, scales as $1/n^{\gamma}$ with $\gamma>1$ ($\gamma=1.83$ for $U=4$ and increases for larger $U$ reaching $\gamma=2.07$ for $U=8$ as shown in Fig. 1 (b). To construct Fig. 1(b), we used DMRG to solve OHK with $n$ as large as 40. Such an orbital number exceeds the limitations of ED (Supplemental Figs. S2 and S3 contain the full data sets). Hence, the calculations in Fig. 1(b) directly parallel the Hubbard-DMRG study in Fig. 3 in the supplement which shows a $1/L$, $L$ the chain length, convergence as is expected for a purely variational technique. In direct contrast lies the OHK convergence of $1/n^{2}$, significantly faster than the $1/L$ in DMRG. Our ability to solve OHK at least in 1d by DMRG reinforces that we are not limited to compute the energy in OHK with ED. To reinforce the simplicity of the method, all calculations were performed on a personal computer. This example demonstrates that OHK can combine with state-of-the-art numerical methods as well as data-driven techniques to predict the $n\rightarrow N_{\rm sites}$ limit with a convergence (see Fig.  S3) that exceeds even the gold-standard, DMRG. More examples can be found in supplemental Fig. S3. We also compare (supplemental Fig. S4) the charge gap, $\Delta$, in both OHK-ED and Hubbard-DMRG. $\Delta$ vanishes in OHK (as it is in the thermodynamic limit) when $U\rightarrow 0$ whereas it is finite in Hubbard-DMRG as limited by the finite system size.
Figure 4: (a) Density of states representing Mott transition at half-filling of the 4-orbital ($2\times 2$) HK model at $\beta=200/t$. (b) Spin susceptibility of half-filled OHK model showing a peak at the antiferromagnetic ordering wave vector ($\pi,\pi)$ and (c) Spectral function $A({\rm k},\omega)$ at half-filling with $U=8$ and zero temperature with a boardening factor $0.2$.
Since we have established the convergence could be as rapid as $1/n^{2}$, we apply the OHK to 2d because unlike Hubbard-DMRG, the OHK model is easily extendable to higher dimensions. All that is necessary is to arrange the $n$ orbitals in a cluster having the symmetry of the underlying lattice. For $n=4$, implemented here as a $2\times 2$ cluster, the matrix $g({\bf k})$, representing the hopping connectivities in this $4-$orbital case is given by
$\left(\begin{array}[]{cccc}0&\varepsilon_{tx}&\varepsilon_{ty}&\varepsilon_{t^{\prime}}\\ \varepsilon_{tx}&0&\varepsilon_{t^{\prime}}&\varepsilon_{ty}\\ \varepsilon_{ty}&\varepsilon_{t^{\prime}}&0&\varepsilon_{tx}\\ \varepsilon_{t^{\prime}}&\varepsilon_{ty}&\varepsilon_{tx}&0\end{array}\right),$ (7)
with $\varepsilon_{tx}=-2t\cos k_{x}$, $\varepsilon_{ty}=-2t\cos k_{y}$ and $\varepsilon_{t^{\prime}}=-4t^{\prime}\cos k_{x}\cos k_{y}$. This matrix is identical in form to that of a $2\times 2$ Hubbard cluster with twisted boundary conditions (TBC) Huang (2020) where $\alpha$ is replaced with the site index $i$ and $k_{x}=\theta_{x}$ and $k_{y}=\theta_{y}$ where $\theta_{x}$ and $\theta_{y}$ define the TBC. With OHK, however, the twist has a physical interpretation as the crystal momentum, thereby placing OHK in the thermodynamic limit. We compute the density of states for a square lattice ($t^{\prime}=0$) and find that any non-zero $U$ is sufficient to eliminate the metallic state producing a vanishing of the density of states at zero energy as shown in Fig. 4a. Regarding the mechanism of the gap, it cannot be associated with antiferromagnetic order for two reasons. First, AF ordering is prohibited by the Mermin-Wagner theoremMermin and Wagner (1966) at finite temperature in 2d. Also, the charge gap forms at temperatures on the order of $U$, well above the AF scale of $t^{2}/U$. Nonetheless, we do find that short-range AF correlations $S(q=(\pi,\pi))=\langle\hat{S}^{z}(q)\hat{S}^{z}(0)\rangle$ are enhanced by $U$ which increase and ultimately saturate as $T$ decreases, all depicted in Fig. 4b. For the half-filled OHK with $n>1$, AF correlations dominate (see supplemental Fig. S8), a qualitative improvement from the band HK model which shows only a diverging ferromagnetic susceptibility. The residual AF correlation in the doped cases (see supplemental Fig. S9) are essential for capturing the underlying intricate physics. The second reason why the Mott gap cannot be atrtributed to spin physics is that even in the presence of a nearest-neighbour hopping $t^{\prime}$, the gap vanishes entirely for $U<U_{c}$ (see supplemental Fig. S10). However, $S(\bf q=(\pi,\pi))$ does not vanish. Consequently, the spin and charge physics are decoupled, an essential feature of Mott physics. Our work here is consistent with the conclusions from quantum cluster methods on the Hubbard modelMaier et al. (2005). Also of note is the narrowing of the bands which is completely absent in the band HK model. We then compute the spectral function $A({\rm k},\omega)$ at $U=8$, shown in Fig. 4c. The dispersion of the leading excitations are in quantitative agreement with cluster pertubation theory for the 2d square Hubbard modelSeki and Yunoki (2016); Huang et al. (2022a) (see supplemental Fig. S11 for the 8-orbital HK result capturing more subtle features). Consequently, OHK with only 4 orbitals already quantitatively describes the known features of the Mott transition.
Given that the kinetic and potential terms do not commute, the number of double occupied sites plays a crucial role in all the ground state properties. To this end, we compute the double occupancy defined as
$D_{n}=\frac{1}{N}\sum_{{\bf k},\alpha}\langle n_{{\bf k}\alpha\uparrow}n_{{\bf k}\alpha\downarrow}\rangle$ (8)
which directly related to the interaction energy. Consider the stark difference between the non-interacting limits of the band and orbital HK models in terms of double occupancy. Regardless of the model, in the non-interacting limit, the expression for $D_{n}$ factorizes
$D^{\rm non-int}_{n}=\frac{1}{N}\sum_{{\bf k},\alpha}\langle n_{{\bf k}\alpha\uparrow}\rangle\langle n_{{\bf k}\alpha\downarrow}\rangle.$ (9)
For the band HK model ($n=1$), half the BZ is doubly occupied. That is, $\langle n_{{\bf k}\alpha\sigma}\rangle=1$ if occupied and zero otherwise. Consequently, $D^{\rm non-int}_{1}=1/2$. On the other hand, for the orbital model, $\langle n_{{\bf k}\alpha\sigma}\rangle=\langle n_{{\bf k}\beta\sigma}\rangle$ for all $\alpha\neq\beta$ by rotational symmetry. In addition, $\langle n_{{\bf k}\alpha\sigma}\rangle=1/2$ in the reduced Brillouin zone is not unity as in the band HK model. As a result, we find that for the orbital HK model, $D^{\rm non-int}_{n>1}=1/4$. Note that this result is identical to the non-interacting limit of the Hubbard model and underscores how drastically band HK differs from its orbital counterpart. Fig. 5a for $D_{n}$ shows this dramatic difference. In band HK, $D$ decreases steadily from $1/2$ and vanishes for $U>W$. However, in orbital HK model, $D_{n}$, starting at $1/4$, just tapers asymptotically as $U$ increases. Note also the rapid convergence between the 2 and 4-orbital cases which matches closely with state-of-the-art auxiliary-field quantum Monte Carlo simulations (AFQMC)Xu et al. (2024) and finite-temperature QMC (FTQMC) White et al. (1989) on the Hubbard model. The inset also shows the double occupancy of the 4-orbital HK model at $1/8$-hole-doping ($x=0.125$) and its benchmark with QMC simulationsXu et al. (2024) (see supplement for explanation of the slight mismatch at $U=0$ due to definition). The agreement underscores that orbital HK converges rapidly to Hubbard physics. Additionally, the computational cost for solving the 4-orbital HK model is considerably lower, both in terms of algorithm development (40 lines of Mathematica) and actual code execution, when compared to QMC or cluster-type DMFT simulations on the Hubbard model. Most crucially, the orbital HK model is more amenable to analytical calculations, an aspect we intend to explore deeper in future studiesTenkila et al. (2024).
The fact that the double occupancy only vanishes asymptotically as $U$ increases in the orbital HK model implies the presence of dynamical mixing between the upper and lower Mott sub-bands. This arises entirely from the non-commutativityHarris and Lange (1967); Meinders et al. (1993); Eskes et al. (1991) of the kinetic and potential energy and appears as $t/U$ corrections to the low-energy spectral weight (LESW). In the atomic limit, the LESW is strictly $2x$ (where $x$ is the doping level) because each hole can be occupied by a spin-up or spin-down electron. Any dynamical corrections to this necessarily generates double occupancy and hence increase the LESW strictly defined as
[…]
The $\omega_{g}$ locates at the spectral gap in the DOS (see examples in supplemental Fig. S12). In Fig. 5b, we compare the LESW of the 4-orbital HK model with Fig. 3 of Ref. [Meinders et al., 1993]. There is no qualitative difference with ED on the (1d) Hubbard modelMeinders et al. (1993); Eskes et al. (1991). Both increase faster than $2x$. The semiconductor (dashed line) and Fermi liquid (dashed-dotted) results are shown for comparison. As the occupied part of the lower band has a weight $1-x$, the total weight in the lower band now exceeds $1+x$. As only $1+x$ electrons can occupy the lower band, dynamical spectral weight transfer (DSWT) defined as $\Lambda(x)-2x$ and plotted in Fig. 5c impliesPhillips (2010) that the spectral weight in the lower band cannot be exhausted by counting electrons alone. Fig. 5c is quantitatively in agreement with the Hubbard model (Fig. 4 of Ref. [Meinders et al., 1993]), the only difference being the maximum value which for 4-orbital HK is $0.2$ whereas it is $0.23$ for the (1d) Hubbard result. Consequently, that the LESW in the orbital HK model increases faster than $2x$ is a profoundly non-trivial result as no analytically solvable model has ever been formulated to capture this feature of the Hubbard model.
Figure 5: (a) double occupancy at half-filling ($x=0$) as a function of $U$ for the orbital HK model with various $n$. The green and gray dots are from auxiliary-field quantum Monte Carlo (AFQMC)Xu et al. (2024) and finite-temperature QMC (FTQMC) simulationsWhite et al. (1989) on the 2D Hubbard model. The inset of panel (a) compares the 4-orbital HK and QMC resultsXu et al. (2024) at $1/8$-hole-doping ($x=0.125$). LESW and DSWT in the exact solution of the 4-orbital HK model ($U=10$) for the hopping parameters shown in panels (b) and (c) respectively at $\beta=30$. The solid line shown with slope $2x$ is the band HK or atomic Hubbard result. The dashed and dotted lines depict the semiconductor and Fermi liquid results respectively. Note, there is no qualitative difference with the ED results for the Hubbard modelMeinders et al. (1993); Eskes et al. (1991). As $t/U$ increases, so does the DSWT.
For any $n<N_{\rm sites}$, the orbital HK model corresponds to a $N_{\rm sites}/n$ copies of $n$-site Hubbard clusters all connected by TBC. The matrix that determines the TBC is purely local as it is determined by the dynamics. Hence, orbital HK is purely local in the rBZn regardless of the number of orbitals used, even $n=1$ corresponding to band HK whose locality is illustrated in Fig. 2. Theorem: The orbital HK model is equivalent to covering the Brillouin zone with $N_{\rm sites}/n$ Hubbard clusters each containing $n$-sites all connected via TBC. Clearly, the $N_{\rm sites}$-cluster Hubbard model results when $n=N_{\rm sites}$. Loosely, the $N_{\rm sites}/n$ identical copies throughout the Brillouin zone can be thought of as replica copies.
Aside from DSWT, OHK also exhibits a pseudogap in the density of states. Shown in Fig. 6 is the density of states for the 4-orbital HK model with (Fig. 6(a,c)) and without (Fig. 6(b,d)) the next-nearest-neighbor hopping, $t^{\prime}$. Fig. 6(a) shows a suppression of the density of states at the chemical potential ($\rho(\omega=0)\lesssim 0.1$) in the underdoped region with $t^{\prime}=-0.25$ when $U\geq W$ ($W=8t$ is the bare bandwidth), indicative of a pseudogap in the absence of superconducting order. The complete tracking of the density of states at zero frequency in Fig. 6(c) displays the trend that as $U$ increases, the pseudo-gap region appears and extends to higher hole-doped density. In contrast, when $t^{\prime}=0$, the pseudo-gap suppression does not arise until $U\gg W$, as shown in Fig. 6(b,d). These observations are consistent with a recent studyHuang (2020) on a $2\times 2$ Hubbard cluster with TBC.
Figure 6: Density of states at varying hole-doped densities with (a) and without (b) the next-nearest-neighbor hopping $t^{\prime}$ for the 4-orbital HK model ($U/t=10,\beta=30/t$). Panel (a) and (b) share the same legend. Only for $t^{\prime}\neq 0$ is $\rho(\omega)$ suppressed at zero frequency in the under-doped region, thereby indicating a pseudogap. Panel (c) and (d) show $\rho(\omega=0)$ with $t^{\prime}=-0.25$ and 00 respectively, as a function of hole-doped density under various $U$.
Finally, we compute the heat capacity of the 4-orbital HK model at half-filling shown in Fig. 7. Most noticeable is the two-peak structure at $U>W$ representing a demarcation of the charge and spin degrees of freedom into high and low-temperature regimes, respectively, consistent with QMC simulations on the Hubbard model Duffy and Moreo (1997); Wang et al. (2022). Also at half-filling, we find a near-crossing of the heat capacity curves as a function of $U$ at a temperature intermediate between the spin and charge excitations. The Maxwell relations governing the entropy dictateVollhardt (1997) such a crossing. Since there is no sign-problem restriction, we can access the low-temperature heat capacity exactly. As we show in Fig. 7(b), the low-temperature heat capacity data follow a power-law increase detailed in the supplement. While such algebraic growth might seem counterintuitive for a Mott insulator, the first-excited state is charge neutralMai et al. (2024); Zhao et al. (2023b) as shown in the supplement. Such charge-neutral excitations determine the low-T behavior of the specific heat and are consistent with the absence of long-range magnetic order.
Figure 7: Heat capacity of the half-filled 4-orbital HK model with various $U$ at (a) high and (b) low temperatures. Both panels share the same legend.
IV Scaling Argument
The hidden power of the OHK model lies in its convergence properties in higher dimensions. This can be established entirely from the number of degrees of freedom that the orbital number introduces at each $k-$point. The quadratic dependence of the Hamiltonian on number operators implies that in 1d, there $n^{2}$ pairs of interactions at each k-point, hence the scaling $1/n^{2}$. In d-dimensions, the number of degrees is now $n^{2d}$ for $n^{d}$-orbital HK. Consequently, the convergence to Hubbard, guaranteed by Eq. (LABEL:ninfty), should scale as $1/n^{2d}$. This explains why so few orbitals are needed in 2d relative to 1d to match existing results. In the extreme limit of $d=\infty$, we find that this convergence factor vanishes for all $n>1$. Hence, $n=1$ or simple band HK is exact in $d=\infty$. This further explains why the spectral functions of band HK have a central peak as the Mott insulator is approached (see Fig. (2)) as in DMFTGeorges et al. (1996). Note focusing on the operator content is valid only if a fixed point exists. We have shown that HK represents a fixed pointZhao et al. (2023a); Huang et al. (2022b) as it contains the most relevant interaction with scaling dimension $-2$ thereby breaking the $Z_{2}$ symmetryZhao et al. (2023a); Huang et al. (2022b) at the Fermi surface of a Fermi liquid. Adding orbitals to each ${\bf k}$-point does not lead to further momentum integrations as such decorations (new degrees of freedom) of each ${\bf k}$-point just appear as discrete summations. Hence, the convergence to the fixed point should just be determined by the total number of degrees freedom at each k-point, namely $n^{2d}$. Consequently, the fixed pointZhao et al. (2023a); Huang et al. (2022b) controls Mott physics.
V Concluding Remarks:
The connection between OHK to Hubbard physics can also be understood in terms of momentum scattering. In the band HK model, the interaction is local in $k$ space and only zero momentum scattering/transfer obtains, while in the Hubbard model, the scattering momenta include all $q$ in the BZ. As we add orbitals to the band HK model, we effectively add additional momentum scattering. For example, the $2$-orbital HK includes $q=(\pi,\pi)$, and $4$-orbital HK includes $q=(\pi,\pi),(0,\pi),(\pi,0)$. A rigorous derivation to show this can be found in the supplement. That few-orbital HK already captures quantatively Hubbard physics supports the flow to Hubbard through a hierachy of momentum scattering.
The rapid convergence to Hubbard points to a new hybrid approach for OHK in higher dimensions: use DMRG for one of the directions and OHK on the others. Because of the rapid convergence of OHK, such an approach would essentially be exact. Consequently, OHK opens up a fundamentally new approach for simulations on strongly correlated electron matter as a result of the fixed pointZhao et al. (2023a). Moreover, given the consistent evolution of physical quantities with increasing orbital number, it is more amenable to machine learning technques than are other methods. This consistent evolution allows machine learning algorithms to detect patterns and make predictions more effectively, which could facilitate the discovery of new phases and transitions in these systems by leveraging the power of data-driven approaches.
. Acknowledgements We thank Barry Bradlyn for helpful comments and Shiwei Zhang for sending us his double occupancy data that is plotted in Fig. 5a (green dots) and George Sawatzky for a helpful e-mail exchange. This work was supported by the Center for Quantum Sensing and Quantum Materials, a DOE Energy Frontier Research Center, grant DE-SC0021238 (P. M. and P. W. P.). PWP also acknowledges NSF DMR-2111379 for partial funding of the HK work which led to these results.